# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v1

@@ V-B Upper Level: Optimal Reward Decision Equilibrium Analysis
There exists a unique Nash equilibrium (NE) [32] in the multi-MSP rewarding game $\Omega$.
@@
We define the Hessian matrix of $\Psi_{n}$ with respect to $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$. The matrix $\bm{\Lambda}_{n}=diag\left(\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}^{2}},\ldots,\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}\right)$ and the second-order partial derivative matrix $\bm{H}_{n}$ is expressed by
@@
$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime},\forall m\in\mathcal{M},$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}},$ (33)
@@
To obtain the values of $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$, we need to obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$, $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$. From (24), we can obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, the steps of which are shown as follows. The first-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-2(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{3}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}=0$. Consequently, we can obtain
by substituting (34) into (35), we have
Likewise, we derive $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ with the following steps based on (26). The first-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-2(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}=0$. Accordingly, we have
by substituting (37) into (38), we have
Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40). Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$. For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (32) and obtain the following result as shown in (41). Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium solution to the multi-MSP rewarding game $\Omega$ has been proved.
$V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}\left(\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
$V_{mn}^{\prime\prime}=\frac{\partial^{2}V_{mn}}{\partial p_{mn}^{2}}=I_{mn}\left[\frac{-2x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{3}}\cdot(\frac{\partial f_{mn}}{\partial p_{mn}})^{2}+\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}\right.$
$\left.-\frac{2b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}log(1/\theta_{m}),a_{2}=log_{2}(1+\varsigma_{mn}).$
1
2Input: Action threshold $[p_{min},p_{max}]$, learning rate;
3 Output: The reward decisions $\bm{p}$ of MSPs, resource allocation $[\bm{f},\bm{B}]$ of MUs;
4 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$ and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ ;
@@ 3 Output: The reward decisions $\bm{p}$ of MSPs, resource allocation $[\bm{f},\bm{B}]$ of MUs;
6 for  Episode = $1,2,\ldots$ do
7      for Epoch $k=1,\ldots,|\mathcal{D}|$ do
8            foreach MSP agent $n\in\mathcal{N}$ do
@@ 9                  Observe space $s_{n}(k)$;
10                   Choose price $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$ by sampling from $\pi_{\alpha_{n}}(s_{n}(k))$ ;
11             end foreach
12            All MSP agents do:
13             Perform a joint reward $\bm{p}(k)=\left\{\bm{p}_{1}(k),\ldots,\bm{p}_{n}(k)\right\}\in\mathcal{A}$;
14             Broadcast reward decisions;
@@ 17             Store $s_{n}(k)=[V_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),V_{n}(k+1)]$ ;
19      Update actor network $\bm{\alpha}$ and critic network $\bm{\beta}$ for each MSP agent as [33] ;
20       Clear up relay buffer $\mathcal{D}$ ;
@@
22# FL training guided by trading results within $T$:
@@
25             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to locally train the task of MSP $n$, and subsequently upload the model to the MSP $n$.
@@
27      Aggregate and issue an updated global model;
@@ 28       Supply the updated model to the AR engine;
We verify the existence and uniqueness of the equilibrium solution for EPEC through Theorem 1 and Theorem 2. However, balancing multiple and conflicting interests, i.e., the utilities of MSPs and MUs, inevitably complicates the optimization problem. To find a solution that optimizes the profits of MSPs while taking into account the utilities of MUs, we develop a reinforcement learning-based strategy in the following section.
@@ VI Dynamic Reward Strategy for MSPs
According to the EPEC proved above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical difficulties: (i) under the time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is non-linear with complicated form and the reward decisions of MSPs are tightly coupled. (ii) MSPs must know private information about MUs (e.g., cost factors) as a prior condition for determining their rewards, which violates the privacy of MUs. To overcome the above difficulties, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [34] to adapt to dynamic channel conditions, where each MSP acts as an individual agent to make intelligent reward decisions.
@@ VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
In our work, each MSP agent operates the DRL-based dynamic reward algorithm in a fully distributed manner. Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 5-19) and the guidance for the FL process (Lines 22-27).
Trading process: When the game starts, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, essential parameters and inference loss for trading (Line 3). During each game, each agent feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 8-9). Then, all the agents generate reward decisions and broadcast them to the MUs (Lines 12-13). After completing these steps, the MSP agents enter the interaction phase with the environment, receive the best feedbacks from the MUs, calculate their utilities, and store information into episode buffer $\mathcal{D}$ (Lines 14-16). The experience is collected through the above interactive process (Lines 8-16) until the buffer $\mathcal{D}$ is full. Then, the critic-network $\bm{\beta}_{n}$ evaluates the actor-network $\bm{\alpha}_{n}$ based on the experience and in turn adjusts it (Line 18). The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase. MUs follow the resource allocation scheme for local training and model uploading (Line 24). Subsequently, the MSP is responsible for aggregating and updating the global model (Line 26). Finally, the updated model is provided for use by the AR engine (Line 27). This process iterates multiple times until the total FL time $T$ is reached. During real-time interactions, the AR service of the MSP can be enhanced in terms of immersion and service quality by the updated model.
@@ VII-B Benchmark Schemes
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [25], the MU allocates computation and communication resources based on the amount of data involved in local training.
@@
  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computation and communication resources to reflect the heterogeneity of devices.
@@
90% Accuracy 50% mAP 70% mAP50
@@
immersion-aware 15.23s 13.09s 7.59s 513.16s 581.09s 623.89s 418.28s 447.69s 451.27s
x_based 17.79s 18.03s 19.80s 812.53s 898.59s 1061.50s 596.57s 679.91s 890.47s
w_based 30s+ 24.04s 10.28s 1064.83s 1175.58s 1200s+ 774.09s 953.50s 1013.89s
w_x_based 23.20s 19.08s 12.34s 901.61s 1190s 1194s 596.35s 848.68s 932.73s
fixed 30s+ 30s+ 30s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+
  • The $1200$s+ and $30$s+ indicate that the guidance time for the current trading is exceeded.
@@ VII-D Benchmark Comparison
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computation and communication resources for providing learning models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training. As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace. In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased. Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
@@
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because the potential value $w$ is higher for MSP 3, leading to more rewards being given by MSP 3. As a result, the MUs will allocate more computation and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource. Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the nature of the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Yet, the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them. The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU. Additionally,from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $w$ associated with MSP 2 and MSP 3.
@@
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Meanwhile, Figs. 10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively. The above comparison validates the feasibility of the proposed trading framework and the advantages of IoM-based metric incentives.

### v2

@@ V-B Upper Level: Optimal Reward Equilibrium among MSPs
There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal set of reward decisions $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
@@
We define the Hessian matrix of $\Psi_{n}$ with respect to its reward vector $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$. The matrix $\bm{\Lambda}_{n}=diag\left(\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}^{2}},\ldots,\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}\right)$ and the second-order partial derivative matrix $\bm{H}_{n}$ is expressed by
@@
$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}<0,\forall m\in\mathcal{M},$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}<0,$ (23)
@@
The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$ can be found in Appendix B. We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved. ∎
1 Input: Action threshold $[p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
2 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n\in\mathcal{N}$;
@@ 1 Input: Action threshold $[p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
4 for episode = $1,2,\ldots,\mathcal{T}$ do
5      Concurrently for each MSP agent $n\in\mathcal{N}$:
6       for epoch $k=1,\ldots,|\mathcal{D}_{n}|$ do
@@ 7            Observe space $s_{n}(k)$;
8             Choose price $\bm{p}_{n}(k)\in[p_{min},p_{max}]$ by sampling from its current policy $\pi_{\alpha_{n}}(s_{n}(k))$;
9             Broadcast reward decisions to MUs;
@@ 12             Store transition $e_{n}(k)=[s_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),s_{n}(k+1)]$;
14      Update actor network $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$ for each MSP agent $n$;
15       Clear the episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$;
@@
17Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $[\bm{f},\bm{B}]$ of MUs;
18 # FL training guided by trading results within time $T$:
@@
21             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the model to MSP $n$.
@@
23      Aggregate received local models and issue an updated global model;
@@ 24       Supply the updated model to the AR engine;
Utilizing Theorem 1 and Theorem 2, we prove the existence and uniqueness of the hierarchical equilibrium in the proposed EPEC model. Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards, while each MSP determines a unique optimal pricing to maximize its own utility given the rewards of other MSPs.
Although the backward induction approach theoretically ensures an equilibrium solution to (14), solving it in real-world vehicular metaverse environments can be challenging due to dynamic network conditions and limited system knowledge (e.g., private cost parameters). In the following section, we address these practical issues by introducing a fully distributed reinforcement learning approach, enabling MSPs to adaptively and privately optimize their reward decisions in real time.
@@ VI Dynamic Reward Strategy for MSPs
According to the EPEC proven above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical challenges: (i) In a time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is a non-linear problem with a complicated structure. Additionally, the reward decisions of MSPs are tightly coupled. (ii) MSPs must have prior knowledge of private information about MUs (e.g., cost factors) to determine their rewards, which raises privacy concerns for MUs. To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [35], which adapts to dynamic channel conditions. In this framework, each MSP is modeled as an individual agent that makes intelligent reward decisions.
@@ VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
In our work, each MSP agent operates the DRL-based dynamic reward algorithm in a fully distributed manner. Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 4-17) and the guidance for the FL process (Lines 19-25).
Trading process: At the start of the game, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ (Line 2). Each agent $n$ feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 7-8). Subsequently, each MSP agent $n$ broadcasts its reward decisions, interacts with the environment, receives feedback from the MUs, calculates its utility, and stores the experience in the episode buffer $\mathcal{D}_{n}$ (Lines 9-12). The experience is collected through this interactive process (Lines 7-12) until the buffer $\mathcal{D}_{n}$ is full. Once the buffer is full, the critic network $\bm{\beta}_{n}$ evaluates the actor network $\bm{\alpha}_{n}$ based on the collected experience and updates it accordingly (Line 14). Finally, the episode buffer is cleared in preparation for the next episode (Line 15). The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase. MUs follow the resource allocation scheme for local training and model uploading (Line 21). Then, the MSP is responsible for aggregating and updating the global model (Line 23). Finally, the updated model is provided for use by the AR engine (Line 24). This process iterates multiple times until the total FL time $T$ is reached. During real-time interactions, the MSP’s AR service immersion and quality can be enhanced by utilizing the updated model.
@@ VII-B Benchmark Schemes
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [31], the MU allocates computational and communication resources based on the amount of data involved in local training.
@@
  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computational and communication resources to reflect the heterogeneity of devices.
@@
Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@5050{50}50 (s)
@@
immersion-aware 15.23 13.09 7.59 513.16 581.09 623.89 418.28 447.69 451.27
x_based 17.79 18.03 19.80 812.53 898.59 1061.50 596.57 679.91 890.47
w_based 30+ 24.04 10.28 1064.83 1175.58 1200+ 774.09 953.50 1013.89
w_x_based 23.20 19.08 12.34 901.61 1190 1194 596.35 848.68 932.73
fixed 30+ 30+ 30+ 1200+ 1200+ 1200+ 1200+ 1200+ 1200+
  • Note: $30$+ or $1200$+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
@@ VII-D Benchmark Comparison
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing learning models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training. As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace. In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased. Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
@@
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because the potential value $\omega$ is higher for MSP 3, leading to more rewards being given by MSP 3. As a result, the MUs will allocate more computational and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource. Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Yet, the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them. The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU. Additionally,from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $\omega$ associated with MSP 2 and MSP 3.
@@
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP@5050{50}50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Meanwhile, Figs. 10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively.
The comparison results validate the advantages of the proposed trading framework, which outperforms other schemes by improving IoM by $38.3\%$ and $37.2\%$ and reducing the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.